The complex numbers z1 z2 and z3 satisfying z1- z3/z2 -z3= 1-i(root 3)/2 are vertices of a triangle which is


 


Of arwa 0


eqiulateral


right angled isosceles


obtuse angle isosceles

Grade:11

1 Answers

Sunil Raikwar
10 years ago
(z1-z2)/(z2-z3)= cos(-60)+sin(-60)=e-i60
taking mode on both sides
Absolute value of(z1-z2/z2-z3)=1
Therefore |z1-z2|=|z2-z3|
Also angle between z1-z2 & z2-z3 is 60
So given triangle is equilateral triangle.
Thanks & Regards
Sunil Raikwar
askIITians faculty